Sum the Series:
1×2+2×3+3×4+4×5+⋯+2013×2014
Thank you!

Answer by rothauserc(4718)

This is a quadratic series whose members are given as
2 6 12 20 30 ......4054182
1 2 3 4 5 2013
n^2 1 4 9 16 25 4052169
a member in the geometric sequence is given by
nth term = n^2 +n
now we know that the summation of n^2 = n[n+1][2n+1]/6 and
summation of n is n(n+1)/2
therefore the sum(S) of the series is
S = 2013(2014)(2*2013+1)/6 + 2013(2014)/2
S = 16326190914/6 + 4054182/2
S = 2721031819 + 2027091
S = 2723058910